OpenStudy (anonymous):

Identify the discontinuities, graph, and find the zeros of each rational function. f(x)=2x^2+5x-12/x+4

OpenStudy (anonymous):

Notice that \[ \frac{2 x^2+5 x-12}{x+4}=\frac{(2 x-3) (x+4)}{x+4}=2 x-3 \] What can you conclude?

OpenStudy (anonymous):

It has a removable singularity at x=-4

OpenStudy (anonymous):

It has a zero at x =3/2

OpenStudy (anonymous):

The graph is the line y= 2 x-3